1) A 90.0-N box is pulled 10.0 m along a level surface by a rope. If the rope makes an angle of 20.0° with the surface, and the force exerted through the rope is 75.0 N, the work is done on the box is ___J.




2) A 25.0-N object is held 2.10 m above the ground. The potential energy of the object with respect to the ground is ___J.



3) The kinetic energy of a 20.0-N object is 5.00 × 102 J. The speed of this object is ___m/s.

4) A 10.0-N object is accelerated uniformly from rest at a rate of 2.5 m/s2. The kinetic energy of this object after it has accelerated a distance of 15.0 m is ___J.

1
Expert's answer
2021-07-14T10:16:13-0400

1)


"W=(75)(10)\cos{20}=705\ J"

2)


"U=(25)(2.1)=52.5\ J"

3)


"0.5\frac{20}{9.8}v^2=500\\v=22.1\frac{m}{s}"

4)

"K=\frac{10}{9.8}(2.5)(15)=38.3\ J"